OpenStudy (anonymous):

(3-2i)(4-7i)----this is imaginary numbers. can anyone solve

OpenStudy (saifoo.khan):

http://www.youtube.com/watch?v=AOke3hRIG34

OpenStudy (anonymous):

FOIL, just remember that i^2=-1

OpenStudy (anonymous):

right. my final answer is 26-29i. Nothing else to be done after that?

OpenStudy (anonymous):

your final answer is wrong, try it again

OpenStudy (anonymous):

But no, after you get a+bi, you can't factor any more

OpenStudy (anonymous):

(3-2i)(4-7i)= 12-21i-8i+14i^2= 12-29i+14i^2= 12-29i-14= -2-29i

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

howd you get from that first step to the second????? dockworker

OpenStudy (anonymous):

using the foil method

OpenStudy (anonymous):

foil??